OpenStudy (anonymous):

Determine the zeros of f(x) = x3 – 12x2 + 28x – 9

OpenStudy (ybarrap):

Factoring this function: $$ f(x) = x^3 – 12x^2 + 28x – 9=(x-9) (x^2-3 x+1) $$ Shows that the zeros occur when $$ x-9=0\\ x^2-3 x+1=0 $$ x=9 is one zero. The other two can be found applying the quadratic formula to the second.

OpenStudy (ybarrap):

Make sense?

OpenStudy (nincompoop):

kind of. does it make sense? what does it mean by zeros? why do we look for zeros?

OpenStudy (ybarrap):

The zeros simply mean where does the function cross the x-axis. That is, when does the function equal zero. That's all.

OpenStudy (ybarrap):

When you have a function of power 3 as you do here, then the maximum number of REAL zeros (i.e. that aren't imaginary0, is also 3.

OpenStudy (anonymous):

Sorry, I just came back. How exactly did you get (x-9)(x^2 +3x -1) ?

OpenStudy (ranga):

To find solutions to this cubic equation, start with the rational roots theorem. The possible roots are: \[\pm \frac{ \text{factors of constant term} }{ \text{factors of leading coefficient} }= \pm \frac{ 1,3,9 }{ 1 }\]

OpenStudy (anonymous):

Leading coefficient would be from the one with the highest degree?

OpenStudy (ybarrap):

I don't have a method that works every time, but I play around with factors that work. If I suspect a factor, I divide out that factor to get the other factors. Here, I suspected 9 was a factor.

OpenStudy (anonymous):

Ah.

OpenStudy (ybarrap):

*(x-9) was a factor

OpenStudy (ranga):

yes. Try the following for x values: 1, -1, 3, -3, 9, -9 until one of them makes f(x) = 0. Then you would have found one root. Do synthetic division to find the quotient.

OpenStudy (ybarrap):

@ranga has stated it perfectly

OpenStudy (anonymous):

@ybarrap, why can't it be (x^2 -9)? Therefore, I get -3 hisand +3 out of this

OpenStudy (anonymous):

@ranga Got it.

OpenStudy (anonymous):

But I just need to find the zeroes, not go deeper to find the quotent.

OpenStudy (ybarrap):

Try it, put in x=3 and x=-3 and see if f(x) = 0. If it does not, then they are not zeros.

OpenStudy (ranga):

Here the rational roots theorem gives you 6 possible roots: 1,-1,3,-3,9 and -9. One of them happens to work. Thanks @ybarrap

OpenStudy (ybarrap):

That right, you will have three values of x that make f(x)=0 and there will be no more because you are limited by the power of 3 to a maximum of 3 zeros.

OpenStudy (anonymous):

Got it. Thanks!

OpenStudy (ranga):

alright.

OpenStudy (ybarrap):

yw